# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 1.3. Strategy of the proof
By pigeonholing, we can find $\mathcal{W}\subset A(1,\mathbb{F}_{p}^{2})$ so that #$\#\mathcal{L}_{W}$ are comparable for all $W\in\mathcal{W}$. Moreover, denoting #$\#\mathcal{W}=p^{t_{1}}$, #$\#\mathcal{L}_{W}=p^{t_{2}}$, we may assume $t_{1}+t_{2}=t$.
@@
By pigeonholing on #$\#E_{W}\cap(\xi\times\mathbb{F}_{p})$, we may assume there is a subset $\Xi_{W}\subset W$ so that #$\#E_{W}\cap(\xi\times\mathbb{F}_{p})$ are comparable for $\xi\in\Xi_{W}$. Moreover, denoting #$\#\Xi=p^{s_{1}}$, #$\#E_{W}\cap(\xi\times\mathbb{F}_{p})=p^{s_{2}}$, we may assume
@@
Also, we may assume $s_{1},s_{2}$ are the same for all $W\in\mathcal{W}$ by doing another pigeonholing on $W$. We can also assume a strong condition on $s_{1}$: $s_{1}\geq s$. But we do not expand the discussion here.
@@
Note that $\bigcup_{W\in\mathcal{W}}\Xi_{W}$ is a $(s_{1},t_{1};2,1)$-set. Hence
#$\#E\gtrsim p^{\mathbf{F}(s_{1},t_{1};2,1)+s_{2}}.$
The problem boils down to the following linear programming problem: suppose
(6) $\begin{cases}t_{1}+t_{2}=t,\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};2,1),\\ s_{1}\geq s.\end{cases}$
@@
The proof of the last inequality will be done by brutal force.
1.4. Structure of the paper
In Section 2, we define $\mathbf{F}(s,t;n,k),\mathbf{M}(a,s;n,k)$ and prove some fundamental lemmas. In Section 3, we prove Theorem 6. In Section 4 and 5, we prove Theorem 7. In Section 6, we prove Theorem 9. In Section 7 and 8, we prove Theorem 10.
Acknowledgement. The author would like to thank Larry Guth, Bochen Liu, Hong Wang and Shukun Wu for some helpful communications.
2. Preliminary
2.1. Definition of $\mathbf{F}(s,t;n,k)$ and $\mathbf{M}(a,s;n,k)$
Now, we make clear the mysterious $\mathbf{F}$ and $\mathbf{M}$ that appeared in Theorem 6, 7, 9, 10.
Definition 11 (Furstenberg index).
@@ 2.1. Definition of $\mathbf{F}(s,t;n,k)$ and $\mathbf{M}(a,s;n,k)$
(9) $\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$
@@
(10) $\begin{split}\mathbf{F}(s,t;n,k)&=d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}\\ &=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}.\end{split}$
@@
(11) $\mathbf{F}(s,t;n,k)=s+m+1(=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}).$
Remark 12.
Through this definition, we can calculate $\mathbf{F}(s,t;2,1)=\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}$ for $s\in(0,1],t\in[0,2]$. One sees that Conjecture 4 is equivalent to saying that $\inf_{E:\ (s,t;2,1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;2,1)-\varepsilon}$.
The following is a construction of $(s,t;2,1)$-set, which provides evidence on why Conjecture 4 should be reasonable.
Proposition 13.
@@
(12) #$\#E\lesssim p^{\max\{0,t-1\}}.$
For $0<s\leq 1,0\leq t\leq 2$.  there exists a $(s,t;2,1;\frac{1}{2})$-set $E$ such that
(13) #$\#E\lesssim p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}}.$
@@
When $s=0$, we construct $E=\bigcup_{L\in\mathcal{L}}Y(L)$ as follows. If $t\leq 1$, choose $\mathcal{L}$ to be a set of $p^{t}$ lines passing through $(0,0)$ and $Y(L)=(0,0)$. Then #$\#E=1=p^{0}$. If $t>1$, we choose $p^{t-1}$ points $E$ on $\mathbb{F}_{p}\times\{0\}$, and for each $x\in E$, we choose $p-1$ lines through $x$ that are not parallel to $\mathbb{F}_{p}\times\{0\}$. If the line $L$ is through $x\in E$, then we let $Y(L)=x$. We see that $E$ is a $(0,t;2,1;\frac{1}{2})$-set with cardinality $p^{t-1}$. (Here, $p^{t-1}$ should actually be the integer $\lceil p^{t-1}\rceil$, but we still write it as $p^{t-1}$ for simplicity, thinking of it as an integer. Same convention applies in the whole paper.)
@@
  1. (1)
    If $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is a $(s,t;2,1;\frac{1}{2})$-set, then #$\#E\leq\sum_{L\in\mathcal{L}}\#Y(L)\leq p^{s+t}$.
  2. (2)
    Choose $\mathcal{L}$ to be a set of $\frac{1}{2}p^{t}$ lines that are not parallel to the line $\mathbb{F}_{p}\times\{0\}$. Let $Y(L)=L\cap(\mathbb{F}_{p}\times\{1,2,\dots,p^{s}\})$. Then $E$ is a $(s,t;2,1;\frac{1}{2})$-set with $E=\bigcup_{L\in\mathcal{L}}Y(L)\subset\mathbb{F}_{p}\times\{1,\dots,p^{s}\}$. Therefore, #$\#E\leq p^{s+1}$.
  3. (3)
    It remains to consider the case when $s+t\leq 2,s\leq t$. This example is of Szemerédi-Trotter type. The heuristic is shown in Figure 1, where we have $\sim p^{\frac{t-s}{2}}$ directions and for each direction, there are $\sim p^{\frac{t+s}{2}}$ lines. We also have a rectangle of size $\sim p^{s}\times p^{\frac{s+t}{2}}$. The rectangle intersects each line $L$ at $\sim p^{s}$ points, which is defined to be $Y(L)$. We also see $E=\bigcup_{L}Y(L)$ is contained in the rectangle. Therefore, #$\#E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
    Let =$\mathcal{L}=\{y=ax+b:a=1,\dots,p^{\frac{t-s}{2}},b=1,\dots,p^{\frac{s+t}{2}}\}$. Here, we view $a,b$ as elements in $\mathbb{F}_{p}$, and hence $y=ax+b$ as a line in $\mathbb{F}_{p}^{2}$. For each $L\in\mathcal{L}$, let $Y(L)=\{(x,y)\in L:x=1,\dots,\frac{1}{2}p^{s}\}$. We see that $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is contained in $\{(x,y):x=1,\dots,\frac{1}{2}p^{s},y=1,\dots,2p^{\frac{s+t}{2}}\}$. Therefore, #$\#E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
Figure 1.
@@
Definition 14 (Marstrand index).
Let $1\leq k<n$ be integers. Let $a\in(0,n],s>0$. We make the following definition for $\mathbf{M}(a,s;n,k)$. We write $a=m+\beta,s=l+\gamma$ where $m,l\in\mathbb{N}$ and $\beta,\gamma\in(0,1]$. For convenience, we call such expression of $a,s$ canonical.
@@
(15) $\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$
@@
(16) $\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$
@@
(17) $\mathbf{M}(a,s;n,k)=-\infty.$
Remark 15.
$\mathbf{M}(a,s;n,k)=-\infty$ corresponds to the case when exceptional set is empty. We remark that in the Euclidean setting, the empty set is assumed to have dimension $\infty$- ∞, in order to satisfy a Fubini-type heuristic $\dim(A\times B)=\dim A\times\dim B$.
@@
(18) $\mathbf{M}(a,s;2,1)=\begin{cases}1&s>\min\{1,a\}\\ \max\{0,2s-a\}&a-1<s\leq\min\{1,a\}\\ -\infty&s\leq a-1.\end{cases}$
We discuss a construction of exceptional set.
Figure 2.
Proposition 16.
Let $a\in(0,2]$, $s\in(\frac{a}{2},\min\{1,a\}]$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{2}$ with #$\#A\gtrsim p^{a}$, such that
#$\#E_{s}(A;2,1)\gtrsim p^{2s-a}.$
@@
The heuristic is shown in Figure 2. Our set $A$ is the lattice points in the rectangle of size $\frac{1}{5}p^{a-s}\times\frac{1}{5}p^{s}$. For any slope $k\in[1,p^{2s-a}]$, we are able to find $<p^{s}$ lines with slope $k$ to cover $A$. Hence, this slope-$k$ direction belongs to the exceptional set.
@@
Note that if $(x,y)\in A$ and $|k|\leq\frac{1}{5}p^{2s-a}$, then $|y-kx|\leq\frac{1}{2}p^{s}$, which means there exists $|m|<\frac{1}{2}p^{s}$ such that $(x,y)\in\ell_{k,m}$. Hence, #$\#\pi_{\ell_{k,0}}^{*}(A)<p^{s}$. This means $E_{s}(A;2,1)\supset\{\ell_{k,0}:|k|\leq\frac{1}{5}p^{2s-a}\}$.
We see that #$\#A\gtrsim p^{a}$, and $E_{s}(A;2,1)$ has cardinality $\gtrsim p^{2s-a}$. ∎
2.2. Useful lemmas
We prove several useful lemmas, which will be used in the later proof. We suggest the reader to skip this part in the first reading.
Lemma 17.
@@ 2.2. Useful lemmas
  1. (1)
    #$\#G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$;
  2. (2)
    #$\#A(k,\mathbb{F}_{p}^{n})\sim p^{(k+1)(n-k)}$;
  3. (3)
    Fix integers $l,k$ with $l+k\geq n$. Let $W_{0}\in A(l,\mathbb{F}_{p}^{n})$. Then #$\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~{}is~{}a~{}}(l+k-n)\textup{-plane}\}\gtrsim p^{(k+1)(n-k)}$.
@@
(1) and (2) are fundamental. We prove (3). $V\cap W_{0}$ is a $(l+k-n)$-plane means $V$ intersects $W_{0}$ transversely. (3) agrees with the intuition in $\mathbb{R}^{n}$, since generic $V\in A(k,\mathbb{R}^{n})$ intersect $W_{0}$ transversely. Noting (2), we see that it suffices to prove
(19) #$\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~{}is~{}a~{}}r\textup{-plane}\}\lesssim p^{(k+1)(n-k)-1},$
for $r>l+k-n$. The proof of (19) is as follows. Choose a $r$-plane $L$ in $W_{0}$. There are #$\#G(r,\mathbb{F}_{p}^{l})\sim p^{(r+1)(l-r)}$ choices for $L$. If $V\cap W_{0}=L$, then $V\supset L$. We have #$\leq\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-r,\mathbb{F}_{p}^{n-r})\sim p^{(k-r)(n-k)}$ choices for $V$ such that $V\cap W_{0}=L$. The two estimates give (19). ∎
Lemma 18.
Let $0\leq l\leq n$, $1\leq k,m\leq n$ be integers satisfying $n-k\geq m-l$ and $l\leq k,m$. If $W\in G(m,\mathbb{F}_{p}^{n})$, then
#$\#\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi^{*}_{V}(W))\leq p^{l}\}\sim p^{k(n-k)-(k-l)(m-l)}.$
@@
(20) #$\#\bigcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\sim p^{k(n-k)-(k+j-m)j}.$
@@
(21) $\begin{split}&\leq\sum_{U\in G(j,W)}\#\{V\in G(n-k,n):V\supset U\}=\#G(j,\mathbb{F}_{p}^{m})\#G(n-k-j,\mathbb{F}_{p}^{n-j})\\ &\sim p^{j(m-j)}p^{(n-k-j)k}=p^{k(n-k)-(k+j-m)j}.\end{split}$
@@
(22) $\begin{split}&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\\ \supset&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\setminus\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}\end{split}$
One can check #$\#\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}=\#G(n-k-j,\mathbb{F}_{p}^{n-j})=p^{k(n-k-j)}$. $\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}$ has cardinality
(23) $\begin{split}&\leq\sum_{j+1\leq j^{\prime}\leq n-k}\#G(j^{\prime}-j,\mathbb{F}_{p}^{m-j})\#G(n-k-j^{\prime},n-j^{\prime})\\ &\sim\sum_{j+1\leq j^{\prime}\leq n-k}p^{(j^{\prime}-j)(m-j^{\prime})+(n-k-j^{\prime})k}\\ &=\sum_{j+1\leq j^{\prime}\leq n-k}p^{k(n-k-j)+(j^{\prime}-j)(m-k-j^{\prime})}\\ &\lesssim p^{k(n-k-j)-1}.\end{split}$
@@
#$\#\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{k(n-k-j)}.$
We can bound the left hand side of (20) by
#$\geq\#\bigsqcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{j(m-j)+k(n-k-j)}=p^{k(n-k)-(k+j-m)j}.$
Lemma 19.
@@
It is easy to check $\mathbf{F}(s,t;n,k)\geq s$, so we just need to check $\mathbf{F}(s,t;n,k)\geq s+t-k(n-k)$. We check four cases as in Definition 11.
@@
Lemma 20.
Suppose $(s,t_{1}+t_{2};n,k)$ is admissible, $t_{1},t_{2}\geq 0$. Then
@@
This is equivalent to showing that for fixed $s,n,k$, the map $t\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz with Lipschitz constant $\leq 1$. We sketch the proof. First, it is not hard to show $t\mapsto\mathbf{F}(s,t;n,k)$ is continuous. Since $\mathbf{F}(s,t;n,k)$ is piecewise defined, we check each formula of $\mathbf{F}(s,t;n,k)$. We see the derivative on $t$ is $\leq 1$. ∎
$\mathbf{F}(s,t;n,k)$ is continuous in $t$, but not in $s$. Nevertheless, we can still say something about $s$.
Lemma 21.
@@
We just note that $s\mapsto\mathbf{F}(s,t;n,k)$ is defined piecewise for $s$ in left-open intervals of form $(d,d+1]$ where $d\in\mathbb{N}$. Moreover, $s\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz in each of these left-open intervals. ∎
@@
Lemma 22.
Each pair $(a,s)$ satisfying $a\in(0,n],s>0$ belongs to exactly one of the four types in Definition 14.

### v2

@@ 1.3 Strategy of the proof
By pigeonholing, we can find $\mathcal{W}\subset A(1,\mathbb{F}_{p}^{2})$ so that $\mathcal{L}_{W}$ are comparable for all $W\in\mathcal{W}$. Moreover, denoting $\mathcal{W}=p^{t_{1}}$, $\mathcal{L}_{W}=p^{t_{2}}$, we may assume $t_{1}+t_{2}=t$.
@@
By pigeonholing on $E_{W}\cap(\xi\times\mathbb{F}_{p})$, we may assume that there is a subset $\Xi_{W}\subset W$ such that $E_{W}\cap(\xi\times\mathbb{F}_{p})$ are comparable for $\xi\in\Xi_{W}$. Moreover, denoting $\Xi=p^{s_{1}}$, $E_{W}\cap(\xi\times\mathbb{F}_{p})=p^{s_{2}}$, we may assume
@@
Also, we may assume $s_{1},s_{2}$ are the same for all $W\in\mathcal{W}$ by doing another pigeonholing on $W$. We can also assume a strong condition on $s_{1}$: $s_{1}\geq s$. However, we do not expand the discussion here.
@@
Note that $\bigcup_{W\in\mathcal{W}}\Xi_{W}$ is a $(s_{1},t_{1};2,1)$-set. Hence,
$E\gtrsim p^{\mathbf{F}(s_{1},t_{1};2,1)+s_{2}}.$
The problem boils down to the following linear programming problem: suppose that
$\begin{cases}t_{1}+t_{2}=t,\\ s_{1}+s_{2}\geq\mathbf{F}(s,t_{2};2,1),\\ s_{1}\geq s.\end{cases}$ (6)
@@
1.4 Structure of the paper
In Section 2, we define $\mathbf{F}(s,t;n,k),\mathbf{M}(a,s;n,k)$ and prove some fundamental lemmas. In Section 3, we prove Theorem 1.7. In Sections 4 and 5, we prove Theorem 1.8. In Section 6, we prove Theorem 1.10. In Sections 7 and 8, we prove Theorem 1.11.
2 Preliminary
[…]
Now, we give the definitions of $\mathbf{F}$ and $\mathbf{M}$ that appeared in Theorem 1.7, 1.8, 1.10, 1.11.
Definition 2.1 (Furstenberg index).
@@ 2 Preliminary
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (9)
@@
$\begin{split}\mathbf{F}(s,t;n,k)&=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}\\ &=d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}.\end{split}$ (10)
@@
$\begin{split}\mathbf{F}(s,t;n,k)&=s+m+\min\{\tau,\frac{1}{2}(\sigma+\tau),1\}\\ &=s+m+1.\end{split}$ (11)
Remark 2.2.
In (c) and (d), $\mathbf{F}$ has the same expression, but we separate them for the convenience of later discussions.
Through this definition, we can calculate $\mathbf{F}(s,t;2,1)=\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}$ for $s\in(0,1],t\in[0,2]$. We see that Conjecture 1.5 is equivalent to saying that $\inf_{E:\ (s,t;2,1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;2,1)-\varepsilon}$.
The following is a construction of a $(s,t;2,1)$-set, which provides evidence on why Conjecture 1.5 should be reasonable.
Proposition 2.3.
@@
$E\lesssim p^{\max\{0,t-1\}}.$ (12)
For $0<s\leq 1,0\leq t\leq 2$.  there exists a $(s,t;2,1)$-set $E$ such that
$E\lesssim p^{\min\{s+t,\frac{3}{2}s+\frac{1}{2}t,s+1\}}.$ (13)
@@
When $s=0$, we construct $E=\bigcup_{L\in\mathcal{L}}Y(L)$ as follows. If $t\leq 1$, choose $\mathcal{L}$ to be a set of $p^{t}$ lines passing through $(0,0)$ and $Y(L)=(0,0)$. Then $E=1=p^{0}$. If $t>1$, we choose $p^{t-1}$ points $E$ on $\mathbb{F}_{p}\times\{0\}$, and for each $x\in E$, we choose p1p-1 lines through $x$ that are not parallel to $\mathbb{F}_{p}\times\{0\}$. If the line $L$ is through $x\in E$, then we let $Y(L)=x$. We see that $E$ is a $(0,t;2,1)$-set with cardinality $p^{t-1}$. (Here, $p^{t-1}$ should actually be the integer $\lceil p^{t-1}\rceil$, but we still write it as $p^{t-1}$ for simplicity, thinking of it as an integer. Same convention applies in the whole paper.)
@@
  1. 1.
    If $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is a $(s,t;2,1)$-set, then $E\leq\sum_{L\in\mathcal{L}}\#Y(L)\leq p^{s+t}$.
  2. 2.
    Choose $\mathcal{L}$ to be a set of $\frac{1}{2}p^{t}$ lines that are not parallel to the line $\mathbb{F}_{p}\times\{0\}$. Let $Y(L)=L\cap(\mathbb{F}_{p}\times\{1,2,\dots,p^{s}\})$. Then $E$ is a $(s,t;2,1)$-set with $E=\bigcup_{L\in\mathcal{L}}Y(L)\subset\mathbb{F}_{p}\times\{1,\dots,p^{s}\}$. Therefore, $E\leq p^{s+1}$.
  3. 3.
    It remains to consider the case when $s+t\leq 2,s\leq t$. This example is of Szemerédi-Trotter type. The heuristic is shown in Figure 1, where we have $\sim p^{\frac{t-s}{2}}$ directions and for each direction, there are $\sim p^{\frac{t+s}{2}}$ lines. We also have a rectangle of size $\sim p^{s}\times p^{\frac{s+t}{2}}$. The rectangle intersects each line $L$ at $\sim p^{s}$ points, which is defined to be $Y(L)$. We also see $E=\bigcup_{L}Y(L)$ is contained in the rectangle. Therefore, $E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
    Let $\mathcal{L}=\{y=ax+b:a=1,\dots,p^{\frac{t-s}{2}},b=1,\dots,p^{\frac{s+t}{2}}\}$. Here, we view $a,b$ as elements in $\mathbb{F}_{p}$, and hence $y=ax+b$ as a line in $\mathbb{F}_{p}^{2}$. For each $L\in\mathcal{L}$, let $Y(L)=\{(x,y)\in L:x=1,\dots,\frac{1}{2}p^{s}\}$. We see that $E=\bigcup_{L\in\mathcal{L}}Y(L)$ is contained in $\{(x,y):x=1,\dots,\frac{1}{2}p^{s},y=1,\dots,2p^{\frac{s+t}{2}}\}$. Therefore, $E\lesssim p^{\frac{3}{2}s+\frac{1}{2}t}$.
Figure 1:
@@
Definition 2.4 (Marstrand index).
Let $1\leq k<n$ be integers. Let $a\in(0,n],s>0$. We give the following definition for $\mathbf{M}(a,s;n,k)$. We write $a=m+\beta,s=l+\gamma$ where $m,l\in\mathbb{N}$ and $\beta,\gamma\in(0,1]$. For convenience, we call such an expression of $a,s$ canonical.
@@
$\mathbf{M}(a,s;n,k)=k(n-k)-(m-l)(k-l)+\max\{2\gamma-(\beta+1),0\}.$ (15)
@@
$\mathbf{M}(a,s;n,k)=k(n-k)-(m+1-l)(k-l)+\max\{2\gamma-\beta,0\}.$ (16)
@@
$\mathbf{M}(a,s;n,k)=-\infty.$ (17)
Remark 2.5.
$\mathbf{M}(a,s;n,k)=$-$\infty$ corresponds to the case when the exceptional set is empty. We remark that in the Euclidean setting, it is assumed that the empty set has dimension -$\infty$, to satisfy a Fubini-type heuristic: $\dim(A\times B)=\dim A+\dim B$.
@@
$\mathbf{M}(a,s;2,1)=\begin{cases}1&s>\min\{1,a\}\\ \max\{0,2s-a\}&a-1<s\leq\min\{1,a\}\\ -\infty&s\leq a-1.\end{cases}$ (18)
We discuss a construction of an exceptional set.
Figure 2:
Proposition 2.6.
Let $a\in(0,2]$, $s\in(\frac{a}{2},\min\{1,a\}]$. Then for $p$ large enough, there exists $A\subset\mathbb{F}_{p}^{2}$ with $A\gtrsim p^{a}$, such that
$E_{s}(A;2,1)\gtrsim p^{2s-a}.$
@@
The heuristic is shown in Figure 2. Our set $A$ is the lattice points in the rectangle of size $\frac{1}{5}p^{a-s}\times\frac{1}{5}p^{s}$. For any slope $k\in[1,p^{2s-a}]$, we can find $p^{s}$ lines with slope $k$ to cover $A$. Hence, this slope-$k$ direction belongs to the exceptional set.
@@
Note that if $(x,y)\in A$ and $|k|\leq\frac{1}{5}p^{2s-a}$, then $|y-kx|\leq\frac{1}{2}p^{s}$, which means that there exists $|m|<\frac{1}{2}p^{s}$ such that $(x,y)\in\ell_{k,m}$. Hence, $\pi_{\ell_{k,0}}^{*}(A)<p^{s}$. This means $E_{s}(A;2,1)\supset\{\ell_{k,0}:|k|\leq\frac{1}{5}p^{2s-a}\}$.
We see that $A\gtrsim p^{a}$, and $E_{s}(A;2,1)$ has cardinality $\gtrsim p^{2s-a}$. ∎
2.2 Useful lemmas
We prove several useful lemmas, which will be used later in the proof. We suggest that the reader skip this part in the first reading.
Lemma 2.7.
@@ 2.2 Useful lemmas
  1. 1.
    $G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$;
  2. 2.
    $A(k,\mathbb{F}_{p}^{n})\sim p^{(k+1)(n-k)}$;
  3. 3.
    Fix integers $l,k$ with $l+k\geq n$. Let $W_{0}\in A(l,\mathbb{F}_{p}^{n})$. Then $\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~is~a~}(l+k-n)\textup{-plane}\}\gtrsim p^{(k+1)(n-k)}$.
@@
(1) and (2) are fundamental. We prove (3). $V\cap W_{0}$ is a (l+kn)(l+k-n)-plane means $V$ intersects $W_{0}$ transversely. (3) agrees with intuition in $\mathbb{R}^{n}$, since generic $V\in A(k,\mathbb{R}^{n})$ intersect $W_{0}$ transversely. Noting (2), we see that it suffices to prove
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\cap W_{0}\textup{~is~a~}r\textup{-plane}\}\lesssim p^{(k+1)(n-k)-1},$ (19)
for r>l+knr>l+k-n. The proof of (19) is as follows. Choose a $r$-plane $L$ in $W_{0}$. There are $G(r,\mathbb{F}_{p}^{l})\sim p^{(r+1)(l-r)}$ choices for $L$. If $V\cap W_{0}=L$, then $V\supset L$. We have $\leq\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-r,\mathbb{F}_{p}^{n-r})\sim p^{(k-r)(n-k)}$ choices for $V$ such that $V\cap W_{0}=L$. The two estimates give (19). ∎
Lemma 2.8.
Let $0\leq l\leq n$, $1\leq k,m\leq n$ be integers that satisfy $n-k\geq m-l$ and $l\leq k,m$. If $W\in G(m,\mathbb{F}_{p}^{n})$, then
$\{V\in G(n-k,\mathbb{F}_{p}^{n}):\#(\pi^{*}_{V}(W))\leq p^{l}\}\sim p^{k(n-k)-(k-l)(m-l)}.$
@@
$\bigcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\sim p^{k(n-k)-(k+j-m)j}.$ (20)
@@
$\begin{split}&\leq\sum_{U\in G(j,W)}\#\{V\in G(n-k,n):V\supset U\}=\#G(j,\mathbb{F}_{p}^{m})\#G(n-k-j,\mathbb{F}_{p}^{n-j})\\ &\sim p^{j(m-j)}p^{(n-k-j)k}=p^{k(n-k)-(k+j-m)j}.\end{split}$ (21)
@@
$\begin{split}&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\\ \supset&\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}\setminus\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}\end{split}$ (22)
One can check $\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U\}=\#G(n-k-j,\mathbb{F}_{p}^{n-j})=p^{k(n-k-j)}$. $\bigcup_{j+1\leq j^{\prime}\leq n-k}\bigcup_{U^{\prime}\in G(j^{\prime},W),U^{\prime}\supset U}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\supset U^{\prime}\}$ has cardinality
$\begin{split}&\leq\sum_{j+1\leq j^{\prime}\leq n-k}\#G(j^{\prime}-j,\mathbb{F}_{p}^{m-j})\#G(n-k-j^{\prime},n-j^{\prime})\\ &\sim\sum_{j+1\leq j^{\prime}\leq n-k}p^{(j^{\prime}-j)(m-j^{\prime})+(n-k-j^{\prime})k}\\ &=\sum_{j+1\leq j^{\prime}\leq n-k}p^{k(n-k-j)+(j^{\prime}-j)(m-k-j^{\prime})}\\ &\lesssim p^{k(n-k-j)-1}.\end{split}$ (23)
@@
$\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{k(n-k-j)}.$
We can bound the left-hand side of (20) by
$\geq\#\bigsqcup_{U\in G(j,W)}\{V\in G(n-k,\mathbb{F}_{p}^{n}):V\cap W=U\}\sim p^{j(m-j)+k(n-k-j)}=p^{k(n-k)-(k+j-m)j}.$
Lemma 2.9.
$\mathbf{F}(s,t;n,k)$ is increasing in $s$ and $t$.
Lemma 2.10.
@@
It is easy to check $\mathbf{F}(s,t;n,k)\geq s$, so we just need to check $\mathbf{F}(s,t;n,k)\geq s+t-k(n-k)$. We check four cases as in Definition 2.1.
@@
Lemma 2.11.
$\mathbf{F}(s,t;n,k)$ is $1$-Lipschitz in $t$. More precisely, suppose that $(s,t_{1}+t_{2};n,k)$ is admissible, $t_{1},t_{2}\geq 0$. Then
@@
This is equivalent to showing that for fixed $s,n,k$, the map $t\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz with Lipschitz constant $\leq 1$. We sketch the proof. First, it is not hard to show that $t\mapsto\mathbf{F}(s,t;n,k)$ is continuous. Since $\mathbf{F}(s,t;n,k)$ is piecewise defined, we check each formula of $\mathbf{F}(s,t;n,k)$. We see that the derivative on $t$ is $\leq 1$. ∎
$\mathbf{F}(s,t;n,k)$ is continuous in $t$, but not in $s$. However, we can still say something about $s$.
Lemma 2.12.
@@
We just note that $s\mapsto\mathbf{F}(s,t;n,k)$ is defined piecewise for $s$ in left-open intervals of the form $(d,d+1]$ where $d\in\mathbb{N}$. Moreover, $s\mapsto\mathbf{F}(s,t;n,k)$ is Lipschitz in each of these left-open intervals. ∎
@@
Lemma 2.13.
Each pair $(a,s)$ satisfying $a\in(0,n],s>0$ belongs to exactly one of the four types in Definition 2.4.